HUBBLE SPACE TELESCOPE DAILY REPORT #2185 PERIOD COVERED: 0000Z (UTC) 08/11/98 - 0000Z (UTC) 08/12/98 Daily Status Report as of 224/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Two Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed Two Sets of WF/PC-2/NIC/1/2 7324 (Multispectral Investigation of Dynamics and Cloud Structure on Neptune) The WF/PC-2 and the Near Infrared Camera (Cameras #1 and #2) were used to make multispectral observations of the dynamics and cloud structure of the planet Neptune. This is to investigate the time dependent vertical structure and spatial distribution of its isolated cloud features and zonal bands, the form and stability of its mean circulation, the evolution of known features, and the possible development of major new features such as great dark spots or outbursts. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed Two Sets of STIS/CCD 7635 (STIS CCD Hot Pixel Annealing) The Space Telescope Imaging Spectrograph (CCD) was used to test the hot pixel annealing procedure. The effectiveness of the CCD hot pixel annealing process is assessed by measuring the dark current behavior before and after annealing and by searching for any window contamination effects. The observations were completed as planned, and no problems were reported. 1.5 Completed NIC/2 7226 (Search for Massive Jupiters) The Near Infrared Camera (Camera One, Two and Three) was used to make a corographic search for massive planets around the nearby, young main sequence star GL879. Because of the extreme youth of this object, any low-mass brown dwarf and planetary companions will still be in a higher luminosity phase and thus easily detectable. A dark calibration image was also taken. A guide star acquisition regressed to Fine Lock Backup (3,0,3) at 223/1614Z due to SSLE on FGS #1. The subsequent single FGS #3 acquisition was successful and the TDF was set on time. The attitude as indicated by the prior updates was good. No other anomalies were noted 1.6 Completed STIS/CCD 8026 (Nearby Starbursts and the Evolutionary Link to AGNs) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude search for star bursting regions in nearby galaxies. The observations were completed as planned, and no anomalies were reported. 1.7 Completed NIC/1/2 7416 (Infrared Imaging of the Dust Shell Structures of the Mira Stars o Ceti and R Leo) The Near Infrared Camera (Cameras One and Two) was used to make observations of the Mira Class star HD14386. Due to its mass loss, the Mira-type variable is surrounded by expanding shells composed of gas and dust. The observations were completed as planned, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 16 Successful : 16 A guide star acquisition regressed to Fine Lock Backup (3,0,3) at 223/1614Z due to SSLE on FGS #1. The subsequent single FGS #3 acquisition was successful and the TDF was set on time. The attitude as indicated by the prior updates was good. Scheduled Reacquisitions : 4 Successful : 4 2.2 FHST Updates: Scheduled : 29 Successful : 29 2.3 Operations Notes: IP testing of TDRS-East was successfully supported from 23/1742Z through 1823Z and TDRS-171 from 223/2028Z through 2119Z. Q Channel 32 kbps telemetry was received with no errors for both periods, three NOOPs and twelve FIVEBLKCKs were transmitted in COUNTCHECK Mode and one NOOP and 46 FIVEBLCKs in the NO VERIFY Mode for a total of 7164 commands with no problems. ESTR 2 was used to record engineering data during the IP test (ROP RD-4). SSR Correctable EDAC Errors were cleared two times (ROP RD-4). CCS Strings "A" and "B" were configured to collect a DF-224 SPC memory dump at 223/1430Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH